# Explaining GPT-4’s Schema of Depression Using Machine Behavior Analysis

### v1

@@ 2 Results
[…]
Except for fatigue, all other symptoms that were explicitly mentioned more than 10% of the time —– including the cardinal symptoms and feelings of worthlessness/guilt —– had a widespread influence on estimating the severity of other unmentioned symptoms. The explicit mention of nearly all the items predicted fatigue. However, fatigue was a very weak indicator of other unmentioned symptoms, despite being explicitly mentioned in 25% of the cases. We identified three pairs of symptoms with strong mutual relationships, where the explicit presence of one symptom influenced the implicit estimation of the other, and vice versa. These pairs include anhedonia and depressed mood, fatigue and psychomotor agitation/retardation, and anhedonia and concentration difficulties.
Uncovering GPT-4’s language of depression: GPT-4 generated reliable reasons for diagnosis
From the left bar chart in Figure 4, it is conclusive that GPT-4 identified explicit markers (positive Cohen’s $d$) of depressive symptoms included words that have been known to be associated with depression from prior studies [46, 47], such as affective words (depressed, feeling low, sad, down), verbs (do), and somatic descriptors (energy, tired, sleep). On the other hand, words from non-spans (negative Cohen’s $d$) mostly consisted of functional words and a few nouns (e.g. pandemic, home). All these words on the left bar chart of Figure 4 had a statistically significant association ($p<0.05$; tested for multiple hypotheses using Bonferroni correction).
In the right bar chart, 12 out of the 25 words with positive Cohen’s $d$ had statistically significant Pearson correlation ($r;p<0.05$) with GPT-4’s total estimated PHQ-9. Essays containing ‘not been depressed’ and mentions of ‘happy’ showed a strong negative association with the total score, and phrases like ‘have been depressed’, ‘sleep’, and ‘low’ (as in low energy) had a positive association. From the set of words with negative Cohen’s $d$, only ‘it’ had a significant but weak association ($r=0.09$) with the total estimated PHQ-9. This showed that GPT-4 doesn’t strongly depend on any meaningful linguistic signal outside its cited explicit markers.
@@ 3 Discussion
Receiving professional mental health support is a multi-step, non-linear process that begins with screening and initial assessment. A significant percentage of individuals do not receive timely support due to social barriers and infrastructural limitations within the traditional mental healthcare system [48, 49]. Though effective mental health treatments exist, most individuals with mental disorders do not receive treatment, in part due to shortages of mental healthcare providers [50, 51]. Digital technologies like GPT-4 which possess unprecedented human-comparable skills and may soon be able to practically address many such challenges. With current consumer costs being around \$30/month [52], models such as GPT-4 may be well-positioned to mitigate well-established inhibitors such as high costs, limited mobility, inadequate insurance coverage, and lack of support for parents [15]. Additionally, it can overcome infrastructural limitations, including the need for cultural adaptation and the dearth of mental health clinicians. However, overcoming these challenges does not guarantee precise assessments due to limitations in sensitivity with current measurement tools. To enhance accuracy, these technologies can leverage passive and quasi-passive indicators from daily life [53, 54] to form more precise assessments and aid the delivery of timely care to those in need.
@@
While studies have focused on LLMs’ generations to identify and measure desirable behaviors for mental health support [55, 56], the current study instead sought to inform how the internal structure of depression is represented within an LLM by analyzing its behavior with respect to symptoms. Hence, we tasked GPT-4 with identifying the key symptomatic markers of depression expressed within human essays. These essays contained exhibits of causes (e.g. substance use, sedentary lifestyle) and effects (e.g. tired, exhausted) of depression through direct or synonymous mentions (e.g. “I don’t like doing things anymore” vs “I don’t see the point of getting out of the bed anymore” both suggest lack of interest). From GPT-4’s assessments on language describing depression, we learn how it views the condition to manifest in people, akin to how practicing therapists understand conditions based on the way patients describe it [57, 58].
We evaluated the viability of GPT-4 for assessing depression using two key principles from measurement theory: reliability and validity. We found that GPT-4 displays moderate-high levels of concurrent validity and internal consistency (reliability) with depressive severity from a well-established rating scale measure, PHQ-9. The overall result of GPT-4’s convergence with self-report was consistent with overall associations of GPT-4 [59] and other LLM[33] with standard depression measures, while the item-level results extend convergence to detection of depression symptoms. Further, we found that GPT-4’s generated explanations were generally faithful [60], with little evidence that it relied on linguistic cues outside its explanations of explicit phrases used to estimate PHQ-9 scores. These findings indicate that GPT-4 can be useful at screening patients and identifying symptoms from their language. Language is an information-rich medium compared to rating scales [61] which can provide avenues for making hypotheses about predisposing, precipitating, and perpetuating factors of psychopathology for a given patient while estimating the symptom severity. Thus, the explanatory output of LLM assessment may free up clinicians to prioritize complex case conceptualization and the creation of carefully tailored treatment plans.
[…]
Highly depression-specific symptoms such as anhedonia, depressed mood, and feelings of worthlessness/guilt [22, 62] had both higher base rates of explicit mentions (Table 3) and higher effects on implicit symptoms (Figure 3), explaining the severity estimations when symptoms were unmentioned in the essays. Symptoms with lower levels of specificity to depression [22], like sleep and appetite, were inferrable in the presence of almost any other symptom. These nonspecific symptoms are shared by depression and anxiety-related disorders [22].
@@
Another important consideration in explaining LLMs is the instructions it is given. Given that the information present in the instructions can affect the performance of these models [63, 64, 65, 66], we instructed GPT-4 to estimate symptoms of depression before estimating the severity of overall depression. This provided a means to understand the cohesiveness of the structure of depression, its faithfulness to what it identifies as explicit markers, and varying difficulty levels in estimating the symptoms. This mirrors studies of LLMs in other domains that have demonstrated improved effectiveness when allowed to break a task into compositional steps [67, 68] or when instructed to carry out specific modules that culminate into a task [69, 70].
@@
Using measurement theory to arrive at explanatory schemas can inform how LLMs function, where they are bound to fail, and how to effectively situate them in the mental health support system to maximize overall benefits. There has been increased advocacy for transparency and explainability [71] of LLMs to improve public faith and understand the risks associated with its usage, especially in areas such as mental health support [72]. Our evaluation framework is intended to help clinicians and decision-makers gain insights about model capabilities and pitfalls so that they can determine the capacity at which contemporary models can be used in the current care pipeline and further improve future systems to yield stronger outcomes.
@@
Our work had several limitations. Foremost, the scope of our results is limited to explaining GPT-4’s behaviors when used as a depression assessment tool. Although other modalities of behavior beyond the patient’s verbal behavior can be used for the assessment, we employ language because it intrinsically contains more information, and the advent of LLMs enables us to leverage the power of language to deliver accurate assessments [61, 73] Further exploration is required to assess whether these models are suitable for use in other stages of the mental health treatment pipeline. It is a multi-stage process, encompassing assessment of several other conditions from time to time, and folding in interventions targeted at addressing the causes of patient’s issues. While the nature of the collected data only allowed us to perform a cross-sectional analysis on a small population predominantly from the United Kingdom, we should also evaluate how these models assess mental health for other random samples of the population. Accurate mental health assessments would require taking longitudinal symptom trajectories into account as well [35] to overcome the limitations of cross-sectional evaluations, which may fail to capture the dynamic nature of symptom progression and individual variability over time. Although transformer-based models have been effective in encoding longitudinal representations of behavioral  [74, 75] and heath markers [76], it is vital to understand how LLMs can effectively disentangle both these trajectories over time to assess the mental health severity.
Altogether, our analysis of GPT-4 in assessing depressive symptoms highlights two key insights: (a) moderate to high convergence with self-reported data and (b) an underlying symptom schema that explained its assessments. GPT-4 ’s schema closely matched self-report for major symptoms, but notably downplayed the link with suicidal ideation and overemphasized psychomotor agitation/retardation. This divergence was somewhat mitigated when suicidal ideation was explicitly mentioned, yet the strong overemphasis on psychomotor symptoms led to weaker alignment with both self-reports and expert ratings. Our findings indicate that GPT-4 has the potential to assist clinicians by efficiently assessing patients and prioritizing cases based on depression symptom severity, optimizing time for those in urgent need. Expanding these methods to other mental health conditions and large language models will deepen our understanding of the benefits and limitations of these tools in enhancing mental health care.
4 Methods
Data
Collection. Our data was collected as part of another study [17]. From 1040 participants in total, 258 (24.8%) screened participants reported having been diagnosed with Major Depressive Disorder (MDD) and 259 (24.8%) with Generalized Anxiety Disorder (GAD); and 491 (47.2%) participants were recruited from an unscreened population. 32 (3.2%) respondents were excluded for not answering the control items (see Measures) correctly. Removing participants that failed attention check and incomplete survey responses, the dataset comprised 963 participants, 571 identified as female, 388 as male, 2 preferred not to say, and 2 did not respond. The average age was 33.3 (SD = 11.2; range = 18 – 77) years. Participants reported being from the U.S. (n = 284) and the U.K. (n = 679). Table S1 shows the summary statistics of the self-reported score for each PHQ-9 symptom in our data.
The Swedish National Ethics Review Board deemed the research study (Ethics Application 2020-00730) exempt from requiring ethical approval according to Swedish Law (see §§ 3-4 of the Act [2003:460] on ethical review of research involving humans in Sweden). The research was carried out according to the Declaration of Helsinki. Participants were recruited from the online platform Prolific [77]. The sample comprised a mixture of screened and unscreened respondents.
Measures. Below we only describe the measures used in this study – for all details regarding the data please refer to Gu et al., 2023 [17]. Open-ended question on depression - questions asking participants to describe their depression was adapted from previous research[78] and validated in Gu et al., 2023 [17] (under review). The questions read:
@@ 4 Methods
Following this, participants were administered PHQ-9 [42], a 9 item questionnaire used to assess symptoms of major depressive disorder as defined in the DSM-IV [57]. The rating scale included nine items targeting symptoms such as “Feeling down, depressed, or hopeless”, referring to experiences from the last two weeks. A four-point rating scale is used (0 = Not at all to 3 = Nearly every day) to indicate the experienced severity of each symptom.
@@
Experiment Design
Prompt. Only the text response prompted to describe one’s depression was sent to GPT-4 111versioned 11/06 and referred to as gpt4-1106 in openAI’s GPT directory along with the instruction described in Table 5. The instruction was composed into three parts:
  1. (A)
  2. (B)
  3. (C)
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
For the first two cases, the symptoms were manually corrected to match the abbreviations provided in Step 1 of the instructions. Type iii generations were rectified by manually reformatting the JSON, and the summaries in type iv generations were removed. The symptoms in type v generations were left out of the analysis.
Cleaning and Filtering GPT-4’s Responses
GPT-4 generated responses which were in JSON format as shown in Figure 6 were turned into tabular data by extracting all the scores into individual columns along with a binary variable indicating whether a symptom was found to be explicit by GPT-4. Table S1 shows the descriptive statistics of the individual symptom scores estimated by GPT-4. The excerpts from the user-written text identified by GPT-4 as explicit symptom markers were extracted using simple regular expressions and stored as a CSV.
Expert Judgements
Two psychology experts independently assessed the depression severity from user-written text on a random sample of 209 instances from the same dataset. They were instructed to read the essay to identify excerpts indicative of PHQ-9 symptoms and estimate a severity score for them. If there were no indicative excerpts, they were asked to estimate the symptom severity based on the essay and the estimated explicit symptom scores. The symptoms accompanied by excerpts would pertain to explicit markers as identified by experts.
Analysis
Item Response Theory. In IRT, data from dichotomous or ordinal items are modeled using one or more latent variables with nonlinear link functions, most commonly logistic functions [79]. For this analysis, we used the 1-PL dichotomous logistic IRT model, aka Rasch model [80] to fit item parameters and assess the presence of each of the nine symptoms. This model constrains the slope parameter of the logistic function to 1.0, with the intercept or item location estimated from the responses. The item location parameter indicates where each item stands in relation to assessing depression. We analyzed the rankings of item locations to compare the two sets of scores across all 955 participants.
Interplay Between Explicit and Implicit Symptoms. For each symptom $i$ from the set of all symptoms $P$ (as defined by the PHQ-9), we trained a linear regression model to predict the GPT-4-estimated score of symptom $i$ when it was not explicitly identified. The model was trained to minimize the sum of squared errors, using the scores of all other explicitly identified symptoms $s_{j}^{exp}$, where $j\in P$ and $j\neq i$, as predictors. The linear regression coefficients $\beta$ represent the association between the implicit score of symptom $i$ and the explicit scores of the remaining symptoms, specifically when symptom $i$ was not explicitly identified by GPT-4 in the user text:
In cases where the score for $s_{j}^{exp}$ was not explicit (i.e., not identified by GPT-4), it was replaced by the mean score for that symptom. This process was repeated 500 times using bootstrapped samples of the data. Coefficients whose 90% confidence intervals included 0 were excluded from Figure 3.
Language Of Depression. For each user-written text, 1, 2, and 3 grams usage was computed as a relative frequency for each user-written text split into two groups, ie., tokens that were present in the GPT-4’s excerpt spans and tokens that were left out of the spans. We removed all n-grams used by less than 10 participants and limited them to n-grams that were only composed of alphabets (filtering n-grams that had numbers and special characters). We computed Cohen’s d [81] using n-grams’ relative frequencies when it was found in explicit spans and the rest of the user text. After applying Bonferroni correction, it resulted in 56 tokens associated with explicit indication of symptoms (positive d) and 30 that were implicit (negative d) that were statistically significant ($p<0.05$).
We computed the n-gram (1, 2, and 3) relative frequencies for each user-written text, and created two separate feature sets containing the relative frequencies of positive and negative Cohen’s d respectively. Both features were filtered to n-grams with a Pointwise Mutual Information [82] of at least 4.0, which reduced it to 45 and 25 features respectively. To uncover if GPT-4 estimates rely on any linguistic signals outside the explicit markers it generated, these sets of positive and negative words’ relative frequencies (computed per user) were correlated with the total PHQ-9 score from GPT-4. This analysis was performed using DLATK[83], a Python library for language analysis.
Declarations
[…]
Funding AVG and RLB were supported in part by NIH grant R01-AA028032, VV was supported in part by a DARPA grant #W911NF-20-1-0306, and HAS was supported in part by both the grants. YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290. JCE was supported by Stanford’s Institute for Human-Centered AI. LF was supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence research on LLMs, and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University. ONEK, KK and VCE were supported by FORTE (STY-2022/0007; 2022-01022).
[…]
CoI/Competing Interest ONEK and KK co-founded a start-up that uses computational language assessments to diagnose mental health problems.
[…]
Ethics Approval and Consent to participate This study was deemed exempt from IRB approval. All study participants’ data was anonymized and they consented to share the data for research.
[…]
Data availability The data for this project will be made available through Github. All data including the original study data, GPT-4 responses, cleaned version, as well as the processed features will be open-sourced for replication.
[…]
Materials availability All materials including the manuscript, display Items (tables and figures), code, and data will be available to the public through Github.
[…]
Code availability All programs for this work were coded in Python and will be open-sourced with the data in the same Github repository. This includes programs that accessed GPT-4 to generate responses using OpenAI’s API, cleaning the data, and our analysis.
[…]
Author Contributions AVG, LF, and HAS designed the research; AVG, YKL, and VV built the analytic tools; ONEK and KK were involved in data collection using Prolific; Clinical psychology experts VCE and KK carried out the expert judgements on the depression essays; AVG, VV, YKL, ECS, JCE, KK, ONEK, VCE, HAS, RLB and LF performed the analysis and wrote the manuscript; TD and AVG made all the display items; All authors made the final decision to submit the manuscript.
Supplementary
The distribution of severity scores for self-report and GPT-4’s assessments over 955 samples from Gu et al., 2023 [17] is shown in  Table S1. While, GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms, showing a different distribution at higher values. This shows that PHQ-9 displays an under-estimation bias of severity scores.
Table S1: Distribution of self-report and GPT-4 estimated scores for each item on PHQ-9 (N=955). GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, but its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms.
Item % Scoring 0 % Scoring 1 % Scoring 2 % Scoring 3
Anhedonia 24.3 28.5 32.5 43.4 24.4 26.5 18.8 1.7
Depressed Mood 23.8 25.0 32.1 24.6 22.0 44.0 22.2 6.5
Sleep 25.1 35.2 23.8 54.4 21.1 10.1 30.0 0.4
Fatigue 15.4 29.3 27.2 50.9 22.7 19.4 34.6 0.4
Appetite 30.6 57.2 25.9 40.0 19.1 2.7 24.4 0.1
Worthlessness / Guilt 29.4 43.3 25.4 37.5 23.0 18.3 22.2 0.9
Difficulty concentrating 28.0 36.0 30.9 58.3 24.1 5.7 16.9 0.0
Psychomotor Agi./ Ret. 55.5 53.3 23.2 45.3 15.0 1.4 6.4 0.0
Suicidal ideation 60.3 87.9 21.4 8.8 10.0 2.3 8.3 1.1

### v2

@@ 3 Results
When depressed mood was not explicitly mentioned, it was largely inferred from explicit mentions of anhedonia ($\beta=.60$). Conversely, anhedonia was influenced by depressed mood ($\beta=.49$) and concentration difficulties ($\beta=.24$). The cardinal symptoms of depression were common explicit drivers of nearly all other symptoms with large effect sizes, underscoring their central role in the network of depressive symptoms. This shows that depressed mood and anhedonia are pivotal for accurately estimating the severity of other less specific symptoms of depression, such as sleep disturbances, appetite changes and fatigue, particularly when they are not directly mentioned in the text.
Furthermore, when individuals mentioned sleep disturbances, GPT-4 moderately used this to estimate other somatic symptoms, such as fatigue ($\beta=.18$) and abnormal appetite ($\beta=.24$). However, the reverse relationship wasn’t significant: sleep disturbances were not inferred from the other somatic markers. However, other physical symptoms didn’t strongly predict sleep disturbances. Instead, sleep issues were primarily inferred from depressed mood ($\beta=.26$).
All symptoms except for fatigue that were explicitly mentioned more than 10% of the time —– including the cardinal symptoms and feelings of worthlessness/guilt —– significantly helped GPT-4 to estimate the severity of unmentioned symptoms. The explicit mention of nearly all the items predicted fatigue. However, fatigue was a very weak indicator of other unmentioned symptoms, despite being explicitly mentioned in 1 out of 4 cases. This reflects fatigue’s nonspecific nature, as it occurs broadly across many conditions and isn’t uniquely informative for depression severity. Three pairs of symptoms showed strong two-way relationships, each explicitly mentioned symptom significantly influencing the other’s implicit estimation and vice versa: anhedonia and depressed mood, fatigue and psychomotor agitation/retardation, and anhedonia and concentration difficulties.
@@ 4 Discussion
Receiving professional mental health support is a multi-step, non-linear process that begins with screening and initial assessment. A significant percentage of individuals do not receive timely support due to social barriers and infrastructural limitations within the traditional mental healthcare system [7, 1]. Though effective mental health treatments exist, most individuals with mental disorders do not receive treatment, in part due to shortages of mental healthcare providers [20, 25]. Digital technologies like GPT-4 which possess unprecedented human-comparable skills and may soon be able to practically address many such challenges. With current consumer costs being around \$20/month, models such as GPT-4 may be well-positioned to mitigate well-established inhibitors such as high costs, limited mobility, inadequate insurance coverage, and lack of support for parents [45]. Additionally, it can overcome infrastructural limitations, including the need for cultural adaptation and the dearth of mental health clinicians.
@@
While studies have focused on identifying and measuring desirable behaviors of LLMs for mental health support [4], the current study instead sought to inform how the internal structure of depression is represented within an LLM by analyzing its behavior with respect to symptoms. Hence, we tasked GPT-4 with identifying the key symptomatic markers of depression expressed within human essays. These essays contained exhibits of causes (e.g. substance use, sedentary lifestyle) and effects (e.g. tired, exhausted) of depression through direct or synonymous mentions (e.g. “I don’t like doing things anymore” vs “I don’t see the point of getting out of the bed anymore” both suggest lack of interest). From GPT-4’s assessments on language describing depression, we learn how it views the condition to manifest in people, akin to how practicing therapists understand conditions based on the way patients describe it [10, 44].
We examined the feasibility of using GPT-4 for depression assessment by evaluating two fundamental psychometric properties: reliability and validity. GPT-4 demonstrated strong concurrent validity, indicating its assessments align closely with established measures of depression, such as the PHQ-9 questionnaire. Additionally, GPT-4 showed high internal consistency, suggesting it reliably captures the interconnected nature of depressive symptoms. Our findings align with previous research showing that LLMs [58, 14] generally correlate well with standard depression scales, and we extend these results to symptom-level assessments. Importantly, GPT-4’s cited explicit markers demonstrated both validity and reliability: the linguistic markers it identified aligned with prior findings [9, 46], and were positively associated with its severity estimates, whereas uncited language from depression essays had no meaningful associations. Such language-based assessments offer richer and more detailed insights than traditional rating scales [24], which provide avenues for making hypotheses about predisposing, precipitating, and perpetuating factors of psychopathology enabling a more personalized treatment planning.
Going beyond evaluating GPT-4’s convergence with self-reports and experts, we inferred its schema of symptoms from its behavior in assessing depression. This schema, a representation of the psychological constructs, that was derived from GPT-4’s responses, was depicted as a web of associations between symptoms. We found GPT-4 and self-report shared commonalities in their schema across the first seven PHQ-9 symptoms, while they differed for psychomotor agitation/ retardation and suicidal thoughts.
Specifically, GPT-4 ’s assessment of suicidal thoughts closely matched expert evaluations and explicit self-reports but struggled to infer it when not directly mentioned. This difficulty highlights the challenge of inferring severe but less frequently mentioned symptoms solely from associated emotional states like depressed mood or anhedonia. In contrast, GPT-4 consistently showed lower convergence in assessing psychomotor disturbances, suggesting that these symptoms are less observable through language compared to others. This finding aligns with prior research indicating that objective behavioral signals, such as speech patterns or facial expressions, are more reliable for assessing psychomotor symptoms due to potential recall biases in self-reports [43]. Notably, both suicidal ideation and psychomotor disturbances demonstrated high discriminant power with respect to latent depression severity (2(c)), underscoring their clinical importance and the need to better understand how GPT-4 diverges in modeling these constructs.
GPT-4 ’s approach to inferring symptoms that are not explicitly mentioned aligns with prior research showing that some symptoms are more central and uniquely tied to depression, while others are broader and commonly appear across different mental health conditions [55]. Highly depression-specific symptoms such as anhedonia, depressed mood, and feelings of worthlessness/guilt [55, 17] were frequently mentioned in language (3(a)) and strongly influenced GPT-4’s inference on other symptoms (Figure 4). This pattern suggests these core symptoms serve as foundational indicators in GPT-4 ’s schema of depression, mirroring clinical practice where these cardinal symptoms heavily guide diagnostic reasoning [11]. Conversely, nonspecific symptoms, such as sleep disturbances and appetite changes – common to both depression and anxiety disorders – could be reliably inferred by GPT-4 from the presence of almost any explicitly mentioned symptom.
@@
Another important consideration in explaining LLMs is the instructions it was given. As the information in instructions can affect the model’s response [3], we instructed GPT-4 to estimate symptoms of depression before the overall depression severity. This provided a means to understand the cohesiveness of the structure of depression, its faithfulness to what it identifies as explicit markers, and varying difficulty levels in estimating the symptoms. This mirrors studies of LLMs in other domains that have demonstrated improved effectiveness when allowed to break a task into compositional steps [56] or when instructed to carry out specific modules that culminate into a task [22].
@@
There has been increased advocacy for transparency and explainability [57] to improve public faith and understand the risks it comes with, especially in areas such as mental health support [29]. Using measurement theory to arrive at explanatory schemas can inform how LLMs function, where they are bound to fail, and how to effectively situate them in the mental health support to maximize overall benefits. Our evaluation framework is intended to help clinicians and decision-makers gain insights about model capabilities and pitfalls to determine the capacity at which contemporary models can be used in the care pipeline and further improve future systems to yield stronger outcomes.
@@
Our work had several limitations. Foremost, the scope of our results is limited to explaining GPT-4’s behaviors when depression is conceptualized through the lens of the PHQ-9. While the PHQ-9 is a widely used and validated tool, it reflects a categorical, DSM-based view of depression that may not fully capture the complexity of psychological distress. Future work can expand this approach to data-driven, dimensional models such as HiTOP [26], which may better capture the heterogeneity and complexity of psychopathology across individuals. Although other modalities beyond language (e.g., speech, behavior) can enrich assessment, we focus on language due to its information-rich nature and the ability of LLMs to leverage this medium for making accurate assessments [24, 52]. Additionally, our study was limited to a cross-sectional analysis of a relatively small, predominantly UK-based sample. Broader validation across diverse populations and contexts is necessary. Accurate mental health assessments would require taking longitudinal symptom trajectories into account [15] to overcome the limitations of cross-sectional evaluations, which fails to capture the dynamic nature of symptom progression and individual variability over time. Although transformer-based models have been effective in encoding longitudinal representations of behavioral [59] and health markers [51], it is vital to understand how effective are LLMs in disentangling both these trajectories over time to assess the mental health.
5 Conclusion
This study uncovered the underlying symptom schema within GPT-4 that explains how its arrived at scores having high convergence with experts and standard self-report tests. While GPT-4 ’s schema closely matched self-report for major symptoms, it notably downplayed the link with suicidal ideation and overemphasized psychomotor agitation/retardation. This divergence was somewhat mitigated when suicidal ideation was explicitly mentioned, suggesting it is more conservative when inferring symptoms of suicidality without explicit evidence. On the other hand, its overemphasis on psychomotor symptoms seemed to blanket bias, happening whether it identified evidence or not. With awareness of these discrepancies, the results suggest GPT-4 has the potential to assist clinicians, such as in screening to optimize time for those in urgent need. More widely, this work demonstrates how a machine behavior approach yields explanations of large language models’ processing of psychological concepts.
Ethics Approval and Consent to participate
This study was deemed exempt from IRB approval. All study participants’ data was anonymized and they consented to share the data for research.
Declarations
Funding. AVG and RLB were supported in part by NIH grant R01-AA028032, VV was supported in part by a DARPA grant #W911NF-20-1-0306, and HAS was supported in part by both the grants. YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290. JCE was supported by Stanford’s Institute for Human-Centered AI. LF was supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence research on LLMs, and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University. ONEK, KK and VCE were supported by FORTE (STY-2022/0007; 2022-01022).
CoI/Competing Interest. ONEK and KK co-founded a start-up that uses computational language assessments to diagnose mental health problems.
Preprints
A preprint of this article is available online: arXiv.2411.13800
Data availability
The data for this project will be made available through Github. All data including the original study data, GPT-4 responses, cleaned version, as well as the processed features will be open-sourced for replication. All programs for this work were coded in Python and will be open-sourced with the data in the same Github repository. This includes programs that accessed GPT-4 to generate responses using OpenAI’s API, cleaning the data, and our analysis.
Acknowledgments
The authors extend their gratitude to Daniel Klein, Roman Kotov, Nicholas Eaton, Dina Vivian, Whitney Ringwald, August Nilsson, and Siddharth Manglik for providing feedback on this work.
Supplementary
The distribution of severity scores for self-report and GPT-4’s assessments over 955 samples from Gu et al., 2024 [16] is shown in Table 2. While, GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms, showing a different distribution at higher values. This shows that PHQ-9 displays an under-estimation bias of severity scores.
Table 2: Distribution of self-report and GPT-4 estimated scores for each item on PHQ-9 (N=955). GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, but its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms.
Item % Scoring 0 % Scoring 1 % Scoring 2 % Scoring 3
Anhedonia 24.3 28.5 32.5 43.4 24.4 26.5 18.8 1.7
Depressed Mood 23.8 25.0 32.1 24.6 22.0 44.0 22.2 6.5
Sleep 25.1 35.2 23.8 54.4 21.1 10.1 30.0 0.4
Fatigue 15.4 29.3 27.2 50.9 22.7 19.4 34.6 0.4
Appetite 30.6 57.2 25.9 40.0 19.1 2.7 24.4 0.1
Worthlessness / Guilt 29.4 43.3 25.4 37.5 23.0 18.3 22.2 0.9
Difficulty concentrating 28.0 36.0 30.9 58.3 24.1 5.7 16.9 0.0
Psychomotor Agi./ Ret. 55.5 53.3 23.2 45.3 15.0 1.4 6.4 0.0
Suicidal ideation 60.3 87.9 21.4 8.8 10.0 2.3 8.3 1.1
Figure 5: Prompt given to the study Participants to express their feeling of depression using language.
@@ Acknowledgments
Machine Behavior Analysis of GPT-5
The decrease in correlations of GPT-5 assessments in comparison to GPT-4 (see Table 1) was explained by its Schema of depression described below. Figure 8(c) shows consistently weaker intersymptom relationship compared to self-report. The item location analysis also shows larger differences in discriminative properties of symptoms in GPT-5 (Figure 8(c)).
GPT-5 had a slightly higher rate of symptom identification (Table 9) than GPT-4 for the less observed symptoms (Suicidal Ideation, Psychomotor agitation/retardation, Difficulty Concentration and Appetite). However, its overall convergence with self-report was under both explicit and implicit identification showed a small decrease.
Figure 8: Similarities and differences between GPT-5’s and self-report’s schema of depressive symptoms.
PHQ9 Item Item Location Rank
GPT-5 Self-Report Item location diff
Anhedonia 2 [2, 2] 3 [3, 3] -1
Depressed Mood 1 [1, 1] 2 [1, 2] -1
Sleep 6 [4, 6] 4 [4, 5] 2
Fatigue 3 [3, 3] 1 [1, 2] 2
Appetite 9 [8, 9] 7 [7, 7] 2
Worthlessness / Guilt 4 [4, 6] 5 [4, 6] -1
Difficulty concentrating 5 [4, 6] 6 [4, 6] -1
Psychomotor 8 [8, 9] 8 [8, 8] 0
Suicidal Ideation 7 [7, 7.5] 9 [9, 9] 0
Figure 9: Contrast of explicit & implicit symptom mentions and convergence for GPT-5. Pearson Correlation ($r$) between GPT-5 and self-reported scores for all symptoms under conditions where symptoms were identified to be explicit and implicit (N=955) by GPT-5. The bar chart in % Explicit column represents the rate of explicit mentions according to GPT-5. Higher correlations were observed consistently when symptoms were identified to be explicitly mentioned.
Symptom Explicit $r$ % Explicit vs. % Implicit Implicit $r$
Anhedonia .72 41.21% .52
Depressed Mood .74 88.81% .51
Sleep .53 12.45% .35
Fatigue .57 17.36% .40
Appetite .70 5.65% .20
Worthlessness / Guilt .56 19.56% .28
Concentration .46 11.82% .31
Psychomotor .04 4.60% .11
Suicidal Ideation .70 12.97% .05
Mean .56 23.82% .30